OpenStudy (anonymous):

it is predicted that the population of a particular state will double every 25 years. a) determine the annual and monthly growth rate. express your answer as percents. b)determine the continuous growth rate per year. Express your answer as a percent.

OpenStudy (anonymous):

if you put it as a function and take log at both side it is easy for you to get the annual growth rate or monthly rate or continuous rate~ here is the example of taking the annual growth rate: (1+r)to the power 25th equals 2~ take ln both side you got 25*ln(1+r)=ln(2) and you got r=1-exp(ln2 to the power 1/25); there is something wrong with inserting the equation~ hope you are able to understand what I have said~ and if you need more you can contact me through the message ~